# Complexity Analysis of Normalizing Constant Estimation: from Jarzynski Equality to Annealed Importance Sampling and beyond

### v1

@@ 23       Sample an independent $\xi\sim\operatorname{\mathcal{N}}\left(0,I_{d}\right)$;
where the vector field $(u^{\ell}_{t})_{t\in[0,T_{\ell}]}$ is chosen such that $X_{t}\sim\mu^{\ell}_{t}$ under $\mathbf{P}^{\ell}$, and in particular, the marginal distributions at 00 and $T_{\ell}$ are $\pi_{\theta_{\ell-1}}$ and $\pi_{\theta_{\ell}}$, respectively. By verifying the Fokker-Planck equation, the following PDE needs to be satisfied:
@@
meaning that $(u^{\ell}_{t})_{t\in[0,T_{\ell}]}$ generates $(\mu^{\ell}_{t})_{t\in[0,T_{\ell}]}$. Similar to the proof of JE (Thm. 6), using the relation between metric derivative and continuity equation (Lem. 2), among all vector fields generating $(\mu^{\ell}_{t})_{t\in[0,T_{\ell}]}$, we choose $(u^{\ell}_{t})_{t\in[0,T_{\ell}]}$ to be the a.s.-unique vector field that satisfies $\|u^{\ell}_{t}\|_{L^{2}(\mu^{\ell}_{t})}=|\dot{\mu}^{\ell}|_{t}$ for Lebesgue-a.e. $t\in[0,T_{\ell}]$, which implies (using the chain rule)
@@
By Lem. 12, we can equivalently write $\mathbf{P}^{\ell}$ as the path measure of the following backward SDE:
@@
Recall that $\mathbf{Q}^{\ell}$ is the path measure of Eq. 12 initialized at $X_{0}\sim\pi_{\theta_{\ell}}$, so $X_{t}\sim\pi_{\theta_{\ell}}$ for all $t\in[0,T_{\ell}]$. By Nelson’s relation (Lem. 12), we can equivalently write $\mathbf{Q}^{\ell}$ as the path measure of
@@
The purpose of writing these two path measures in the way of backward SDEs is to avoid the extra term of the KL divergence between the initialization distributions $\pi_{\theta_{\ell-1}}$ and $\pi_{\theta_{\ell}}$ at time 00 when calculating $\operatorname{KL}(\mathbf{P}^{\ell}\|\mathbf{Q}^{\ell})$. To see this, by Girsanov theorem (Lem. 15), the triangle inequality, and the change-of-variable formula, we have
@@
to make the l.h.s. of Eq. 32 $O(\varepsilon^{2})$. Notably, $T$ is the summation over all $T_{\ell}$’s, which has the same order as the total time $T$ for running JE (Eq. 6) in the continuous scenario, in Thm. 5. Plugging this $T_{\ell}$ into the second summation, and noticing that by Eq. 11 and Lem. 33,
@@
we conclude that Eq. 29 implies Eq. 33.
@@
Now consider the other constraint Eq. 28. By chain rule and data-processing inequality,
@@
and ${\mathbf{\widehat{Q}}}^{\ell}$ is the path measure of Eq. 14 initialized at $X_{0}\sim\pi_{\theta_{\ell-1}}$, i.e.,
@@
By Lem. 1, triangle inequality, and the smoothness of $V$, we have
@@
Recall that the potential of $\pi_{\theta}$ is $(\beta+\lambda(\theta))$-smooth. By Lem. 32 and the monotonicity of $\lambda(\cdot)$,
@@
Assume $\max_{\ell\in\left[\!\left[1,M\right]\!\right]}T_{\ell}\lesssim\frac{1}{\beta}$, i.e., Eq. 31. so $\max_{\ell\in\left[\!\left[1,M\right]\!\right]}T_{\ell}(\beta+\lambda(\theta_{\ell-1}))\lesssim 1$, due to $\lambda(\cdot)\leq 2\beta$. We can further simplify the above expression to
@@
So Eq. 31 implies that the r.h.s. of the above equation is $O(1)$. ∎
Finally, we have arrived at the last step of proving Thm. 8, that is to decide the schedule of $\theta_{\ell}$’s.
Define $\vartheta_{\ell}:=1-\theta_{\ell}$, $\ell\in\left[\!\left[0,M\right]\!\right]$. We consider the annealing schedule $\lambda(\theta)=2\beta(1-\theta)^{r}$ for some $1\leq r\lesssim 1$, and to emphasize the dependence on $r$, we use $\mathcal{A}_{r}$ to represent the action of $(\pi_{\theta})_{\theta\in[0,1]}$. The l.h.s. of Eq. 29 is
@@
where the last inequality comes from Lem. 28. So to satisfy Eq. 29, it suffices to ensure
@@
Appendix E Proofs for Sec. 6
E.1 Proof of Prop. 9
@@
The claim of smoothness follows from Guo et al. (2025, Lem. 7). Throughout this proof, let $\phi$ and $\Phi$ denote the p.d.f. and c.d.f. of the standard normal distribution $\operatorname{\mathcal{N}}\left(0,1\right)$, respectively. Unless otherwise specified, the integration ranges are assumed to be (,)($\infty$,$\infty$)( - ∞ , ∞ ).
@@
$=\frac{1}{1+\mathrm{e}^{-\frac{\lambda m^{2}}{2(\lambda+1)}}}\operatorname{\mathcal{N}}\left(x\left|0,\frac{1}{\lambda+1}\right.\right)+\frac{\mathrm{e}^{-\frac{\lambda m^{2}}{2(\lambda+1)}}}{1+\mathrm{e}^{-\frac{\lambda m^{2}}{2(\lambda+1)}}}\operatorname{\mathcal{N}}\left(x\left|\frac{m}{\lambda+1},\frac{1}{\lambda+1}\right.\right).$
@@
Consider change of variable $y=F_{s}(x)$, then $\frac{\mathrm{d}y}{\mathrm{d}x}=f_{s}(x)$. As $x=F_{s}^{-1}(y)$, $(F_{s}^{-1})^{\prime}(y)=\frac{\mathrm{d}x}{\mathrm{d}y}=\frac{1}{f_{s}(x)}$. Taking derivation of $s$ on both sides of the equation $x=F_{s}^{-1}(F_{s}(x))$ yields
00 0 $=\partial_{s}F_{s}^{-1}(F_{s}(x))+(F_{s}^{-1})^{\prime}(F_{s}(x))\partial_{s}F_{s}(x)$
$=\partial_{s}F_{s}^{-1}(y)+\frac{1}{f_{s}(x)}\partial_{s}F_{s}(x).$
@@
$-\partial_{s}F_{s}(x)\gtrsim\frac{m^{2}}{\mathrm{e}^{(1-s)m^{2}/2}}.$
@@
|˙$|\dot{}\underline{\pi}|_{s}^{2}$ $=\int\frac{(\partial_{s}F_{s}(x))^{2}}{f_{s}(x)}\mathrm{d}x\geq\int_{\frac{m}{2}-0.1}^{\frac{m}{2}+0.1}\frac{(\partial_{s}F_{s}(x))^{2}}{f_{s}(x)}\mathrm{d}x$
@@
Since
$|S^{\prime}(\theta)|=\frac{m^{2}r(1-\theta)^{r-1}}{(1+m^{2}(1-\theta)^{r})^{2}}\geq\frac{m^{2}r\left(\frac{1/0.99-1}{m^{2}}\right)^{1-1/r}}{\left(1+m^{2}\left(\frac{1/0.9-1}{m^{2}}\right)\right)^{2}}\gtrsim m^{2/r}\gtrsim 1,\forall\theta\in\Theta,$
the proof is complete. ∎
@@
Define the probability distributions $\underline{\pi}_{s}$ as in the proof of Prop. 9, for some large enough $m\gtrsim 1$. Then, for all $s\in[0.9,0.99]$, we have
@@
Note that $(\pi_{t})_{t\in[0,\infty)}$ satisfies the Fokker-Planck equation $\partial_{t}\pi_{t}=\nabla\cdot\left(\pi_{t}\nabla\log\frac{\pi_{t}}{\gamma}\right)$. Hence, the vector field $\left(v_{t}:=-\nabla\log\frac{\pi_{t}}{\gamma}\right)_{t\in[0,\infty)}$ generates $(\pi_{t})_{t\in[0,\infty)}$, and each $v_{t}$ can be written as a gradient field of a potential function. Thus, by the uniqueness result in Lem. 2, we conclude that
$|\dot{\pi}|^{2}_{t}=\left\|\nabla\log\frac{\pi_{t}}{\gamma}\right\|^{2}_{L^{2}(\pi_{t})}=\operatorname{FI}(\pi_{t}\|\gamma)\leq\mathrm{e}^{-2t}\operatorname{FI}(\pi\|\gamma),$
where $\operatorname{FI}$ is the Fisher divergence, and the last inequality is due to Villani (2003, Eq. 9.34). Finally, using the smoothness of $V$ and Lem. 32, we have
@@
By Nelson’s relation (Lem. 12), $\mathbb{Q}$ is equivalent to the path measure of the following SDE:
$\mathrm{d}X_{t}=X_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}B^{\leftarrow}_{t},~{}t\in[0,T-\delta];~{}X_{T-\delta}\sim\underline{\pi}_{\delta}.$
Leveraging Girsanov theorem (Lem. 15), we know that for a.s. $X\sim\mathbb{Q}^{\dagger}$:
$\log\frac{\mathrm{d}\mathbb{Q}^{\dagger}}{\mathrm{d}\mathbb{Q}}(X)$
$=\log\frac{\phi(X_{0})}{\underline{\pi}_{\delta}(X_{T-\delta})}+(T-\delta)d+\int_{0}^{T-\delta}\left(\|s_{T-t_{-}}(X_{t_{-}})\|^{2}\mathrm{d}t+\sqrt{2}\left\langle s_{T-t_{-}}(X_{t_{-}}),\mathrm{d}B_{t}\right\rangle\right)$
$=\log Z+W(X)+\log\frac{\mathrm{d}\pi}{\mathrm{d}\underline{\pi}_{\delta}}(X_{T-\delta}).$
@@
$Z=\operatorname{\mathbb{E}}_{\mathbb{Q}^{\dagger}(X)}\mathrm{e}^{-W(X)}\frac{\mathrm{d}\underline{\pi}_{\delta}}{\mathrm{d}\pi}(X_{T-\delta}).$
Since $\frac{\widehat{Z}}{Z}=\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{Q}^{\dagger}}(X)\frac{\mathrm{d}\pi}{\mathrm{d}\underline{\pi}_{\delta}}(X_{T-\delta})$, we have
$\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\geq\varepsilon\right)$ $=\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{Q}^{\dagger}}(X)\frac{\mathrm{d}\pi}{\mathrm{d}\underline{\pi}_{\delta}}(X_{T-\delta})-1\right|\geq\varepsilon\right)$
$\leq\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{Q}^{\dagger}}(X)-1\right|\gtrsim\varepsilon\right)+\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\underline{\pi}_{\delta}}(X_{T-\delta})-1\right|\gtrsim\varepsilon\right).$
The inequality is due to the fact that $|ab-1|\geq\varepsilon$ implies $|a-1|\geq\frac{\varepsilon}{3}$ or $|b-1|\geq\frac{\varepsilon}{3}$ for $\varepsilon\in[0,1]$. It suffices to make both terms above $O(1)$. To bound the first term, we use the similar approach as in the proof of Eq. 20 in Thm. 6:
@@
$\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\underline{\pi}_{\delta}}(X_{T-\delta})-1\right|\gtrsim\varepsilon\right)$ $\leq\operatorname{Pr}_{X\sim\mathbb{Q}}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\underline{\pi}_{\delta}}(X_{T-\delta})-1\right|\gtrsim\varepsilon\right)+\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\dagger})$
$\leq\underline{\pi}_{\delta}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\underline{\pi}_{\delta}}-1\right|\gtrsim\varepsilon\right)+\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\dagger})$
$\lesssim\frac{\operatorname{TV}(\underline{\pi}_{\delta},\pi)}{\varepsilon}+\varepsilon.$
Therefore, it suffices to make $\operatorname{TV}(\underline{\pi}_{\delta},\pi)\lesssim\varepsilon$. ∎

### v2

@@ 24   Sample an independent $\xi\sim\operatorname{\mathcal{N}}\left(0,I_{d}\right)$;
where the vector field $(u^{\ell}_{t})_{t\in[0,T_{\ell}]}$ is chosen such that $X_{t}\sim\mu^{\ell}_{t}$ under $\mathbf{P}^{\ell}$, and in particular, the marginal distributions at 0 and $T_{\ell}$ are $\pi_{\theta_{\ell-1}}$ and $\pi_{\theta_{\ell}}$, respectively. By verifying the Fokker-Planck equation, the following PDE needs to be satisfied:
@@
meaning that $(u^{\ell}_{t})_{t\in[0,T_{\ell}]}$ generates $(\mu^{\ell}_{t})_{t\in[0,T_{\ell}]}$. Similar to the proof of JE (Thm.˜2), using the relation between metric derivative and continuity equation (Lem.˜5), among all vector fields generating $(\mu^{\ell}_{t})_{t\in[0,T_{\ell}]}$, we choose $(u^{\ell}_{t})_{t\in[0,T_{\ell}]}$ to be the a.s.-unique vector field that satisfies $\|u^{\ell}_{t}\|_{L^{2}(\mu^{\ell}_{t})}=|\dot{\mu}^{\ell}|_{t}$ for Lebesgue-a.e. $t\in[0,T_{\ell}]$, which implies (using the chain rule)
@@
By Lem.˜3, we can equivalently write $\mathbf{P}^{\ell}$ as the path measure of the following backward SDE:
@@
Recall that $\mathbf{Q}^{\ell}$ is the path measure of ˜8 initialized at $X_{0}\sim\pi_{\theta_{\ell}}$, so $X_{t}\sim\pi_{\theta_{\ell}}$ for all $t\in[0,T_{\ell}]$. By Nelson’s relation (Lem.˜3), we can equivalently write $\mathbf{Q}^{\ell}$ as the path measure of
@@
The purpose of writing these two path measures in the way of backward SDEs is to avoid the extra term of the KL divergence between the initialization distributions $\pi_{\theta_{\ell-1}}$ and $\pi_{\theta_{\ell}}$ at time 0 when calculating $\operatorname{KL}(\mathbf{P}^{\ell}\|\mathbf{Q}^{\ell})$. To see this, by Girsanov theorem (Lem.˜4), the triangle inequality, and the change-of-variable formula, we have
@@
to make the l.h.s. of ˜30 $O(\varepsilon^{2})$. Notably, $T$ is the summation over all $T_{\ell}$’s, which has the same order as the total time $T$ for running JE (˜2) in the continuous scenario, in Thm.˜1. Plugging this $T_{\ell}$ into the second summation, and noticing that by ˜7 and Lem.˜15,
@@
we conclude that ˜27 implies ˜31.
@@
Now consider the other constraint ˜26. By chain rule and data-processing inequality,
@@
and $^{\mathbf{\widehat{Q}}}^{\ell}$ is the path measure of ˜10 initialized at $X_{0}\sim\pi_{\theta_{\ell-1}}$, i.e.,
@@
By Lem.˜2, triangle inequality, and the smoothness of $V$, we have
@@
Recall that the potential of $\pi_{\theta}$ is $(\beta+\lambda(\theta))$-smooth. By Lem.˜14 and the monotonicity of $\lambda(\cdot)$,
@@
Assume $\max_{\ell\in\left[\!\left[1,M\right]\!\right]}T_{\ell}\lesssim\frac{1}{\beta}$, i.e., ˜29. so $\max_{\ell\in\left[\!\left[1,M\right]\!\right]}T_{\ell}(\beta+\lambda(\theta_{\ell-1}))\lesssim 1$, due to $\lambda(\cdot)\leq 2\beta$. We can further simplify the above expression to
@@
So ˜29 implies that the r.h.s. of the above equation is $O(1)$.
Finally, we have arrived at the last step of proving Thm.˜4, that is to decide the schedule of $\theta_{\ell}$’s.
Define $\vartheta_{\ell}:=1-\theta_{\ell}$, $\ell\in\left[\!\left[0,M\right]\!\right]$. We consider the annealing schedule $\lambda(\theta)=2\beta(1-\theta)^{r}$ for some $1\leq r\lesssim 1$, and to emphasize the dependence on $r$, we use $\mathcal{A}_{r}$ to represent the action of $(\pi_{\theta})_{\theta\in[0,1]}$. The l.h.s. of ˜27 is
@@
where the last inequality comes from Lem.˜10. So to satisfy ˜27, it suffices to ensure
@@
Remark 7.
The work Guo et al. (2025) used similar methodologies to prove an $\widetilde{O}\left(\frac{d\beta^{2}\mathcal{A}^{2}}{\varepsilon^{6}}\right)$ oracle complexity for obtaining a sample that is $\varepsilon^{2}$-close in KL divergence to the target distribution. While our assumptions are mostly the same with Guo et al. (2025) except for some insignificant technical ones, and both proofs involve the standard discretization analysis through Girsanov theorem, the improvement of the $\varepsilon$-dependency in Thm.˜4 is due to the fact that Guo et al. (2025) requires $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})\lesssim\varepsilon^{2}$ for sampling, which results in a $\widetilde{\Theta}(\varepsilon^{4})$ step size in Guo et al. (2025), while our proof only needs $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})\lesssim 1$ and $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})\lesssim\varepsilon^{2}$ for normalizing constant estimation, resulting in an improved $\widetilde{\Theta}(\varepsilon^{2})$ step size.
Appendix E Proofs for Sec.˜5
E.1 Proof of Prop.˜1
@@
The claim of smoothness follows from Guo et al. (2025, Lem. 7). A similar approach for proving the lower bound of metric derivative was used independently in Chemseddine et al. (2025, App. B).
Throughout this proof, let $\phi$ and $\Phi$ denote the p.d.f. and c.d.f. of the standard normal distribution $\operatorname{\mathcal{N}}\left(0,1\right)$, respectively. Unless otherwise specified, the integration ranges are assumed to be $(-\infty,\infty)$.
@@
[…]
@@
Consider change of variable $y=F_{s}(x)$, then $\frac{\mathrm{d}y}{\mathrm{d}x}=f_{s}(x)$. As $x=F_{s}^{-1}(y)$, $(F_{s}^{-1})^{\prime}(y)=\frac{\mathrm{d}x}{\mathrm{d}y}=\frac{1}{f_{s}(x)}$. Taking the derivative of $s$ on both sides of the equation $x=F_{s}^{-1}(F_{s}(x))$ yields
$0=\partial_{s}F_{s}^{-1}(F_{s}(x))+(F_{s}^{-1})^{\prime}(F_{s}(x))\partial_{s}F_{s}(x)=\partial_{s}F_{s}^{-1}(y)+\frac{1}{f_{s}(x)}\partial_{s}F_{s}(x).$
@@
$\partial_{s}F_{s}(x)\gtrsim\frac{m^{2}}{\mathrm{e}^{(1-s)m^{2}/2}}.$
@@
$|\dot{}\underline{\pi}|_{s}^{2}$ $=\int\frac{(\partial_{s}F_{s}(x))^{2}}{f_{s}(x)}\mathrm{d}x\geq\int_{\frac{m}{2}-0.1}^{\frac{m}{2}+0.1}\frac{(\partial_{s}F_{s}(x))^{2}}{f_{s}(x)}\mathrm{d}x$
@@
For any $\theta\in\Theta$,
$|S^{\prime}(\theta)|=\frac{m^{2}r(1-\theta)^{r-1}}{(1+m^{2}(1-\theta)^{r})^{2}}\geq\frac{m^{2}r\left(\frac{1/0.99-1}{m^{2}}\right)^{1-1/r}}{\left(1+m^{2}\left(\frac{1/0.9-1}{m^{2}}\right)\right)^{2}}=\frac{m^{2/r}r(1/99)^{1-1/r}}{(1/0.9)^{2}}\gtrsim m^{2/r}\gtrsim 1,$
where in the first $\gtrsim$” we used the inequality $r\left(\frac{1}{99}\right)^{1-\frac{1}{r}}\geq\frac{1}{\mathrm{e}^{4}}$ that holds for all $r\geq 1$. Thus, the proof is complete. ∎
@@
Define the probability distributions $\underline{\pi}_{s}$ as in the proof of Prop.˜1, for some large enough $m\gtrsim 1$. Then, for all $s\in[0.9,0.99]$, we have
@@
Note that $(\overline{\pi}_{t})_{t\in[0,\infty)}$ satisfies the Fokker-Planck equation $\partial_{t}\overline{\pi}_{t}=\nabla\cdot\left(\overline{\pi}_{t}\nabla\log\frac{\overline{\pi}_{t}}{\gamma}\right)$. Hence, the vector field $\left(v_{t}:=-\nabla\log\frac{\overline{\pi}_{t}}{\gamma}\right)_{t\in[0,\infty)}$ generates $(\overline{\pi}_{t})_{t\in[0,\infty)}$, and each $v_{t}$ can be written as a gradient field of a potential function. Thus, by the uniqueness result in Lem.˜5, we conclude that
$|\dot{}\overline{\pi}|^{2}_{t}=\left\|\nabla\log\frac{\overline{\pi}_{t}}{\gamma}\right\|^{2}_{L^{2}(\overline{\pi}_{t})}=\operatorname{FI}(\overline{\pi}_{t}\|\gamma)\leq\mathrm{e}^{-2t}\operatorname{FI}(\pi\|\gamma),$
where $\operatorname{FI}$ is the Fisher divergence, and the last inequality is due to Villani (2003, Eq. 9.34). Finally, using the smoothness of $V$ and Lem.˜14, we have
@@
By Nelson’s relation (Lem.˜3), $\mathbb{Q}$ is equivalent to the path measure of the following SDE:
$\mathrm{d}X_{t}=X_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}B^{\leftarrow}_{t},~t\in[0,T-\delta];~X_{T-\delta}\sim\overline{\pi}_{\delta}.$
Leveraging Girsanov theorem (Lem.˜4), we know that for a.s. $X\sim\mathbb{Q}^{\dagger}$:
$\log\frac{\mathrm{d}\mathbb{Q}^{\dagger}}{\mathrm{d}\mathbb{Q}}(X)$ $=\log\frac{\phi(X_{0})}{\overline{\pi}_{\delta}(X_{T-\delta})}+\frac{1}{2}\int_{0}^{T-\delta}\left(\left\langle X_{t}+2s_{T-t_{-}}(X_{t_{-}}),\mathrm{d}X_{t}\right\rangle-\left\langle X_{t},*\mathrm{d}X_{t}\right\rangle\right)$
$-\frac{1}{4}\int_{0}^{T-\delta}\left(\|X_{t}+2s_{T-t_{-}}(X_{t_{-}})\|^{2}-\|X_{t}\|^{2}\right)\mathrm{d}t.$
Note that for $X\sim\mathbb{Q}^{\dagger}$, $\int_{0}^{T-\delta}\left\langle X_{t},*\mathrm{d}X_{t}\right\rangle=\int_{0}^{T-\delta}\left\langle X_{t},\mathrm{d}X_{t}\right\rangle+[X,X]_{T-\delta}$ and $X,X]_{T-\delta}=\left[\sqrt{2}B,\sqrt{2}B\right]_{T-\delta}=2(T-\delta)d$. Some simple calculations yield
$\log\frac{\mathrm{d}\mathbb{Q}^{\dagger}}{\mathrm{d}\mathbb{Q}}(X)$ $=\log\frac{\phi(X_{0})}{\overline{\pi}_{\delta}(X_{T-\delta})}-(T-\delta)d+\int_{0}^{T-\delta}\left(\|s_{T-t_{-}}(X_{t_{-}})\|^{2}\mathrm{d}t+\sqrt{2}\left\langle s_{T-t_{-}}(X_{t_{-}}),\mathrm{d}B_{t}\right\rangle\right)$
$=\log Z+W(X)+\log\frac{\mathrm{d}\pi}{\mathrm{d}\overline{\pi}_{\delta}}(X_{T-\delta}).$
@@
$Z=\operatorname{\mathbb{E}}_{\mathbb{Q}^{\dagger}(X)}\mathrm{e}^{-W(X)}\frac{\mathrm{d}\overline{\pi}_{\delta}}{\mathrm{d}\pi}(X_{T-\delta})\approx\operatorname{\mathbb{E}}_{\mathbb{Q}^{\dagger}(X)}\mathrm{e}^{-W(X)}=\operatorname{\mathbb{E}}\widehat{Z}$.
Since $\frac{\widehat{Z}}{Z}=\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{Q}^{\dagger}}(X)\frac{\mathrm{d}\pi}{\mathrm{d}\overline{\pi}_{\delta}}(X_{T-\delta})$, we have
$\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\geq\varepsilon\right)$ $=\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{Q}^{\dagger}}(X)\frac{\mathrm{d}\pi}{\mathrm{d}\overline{\pi}_{\delta}}(X_{T-\delta})-1\right|\geq\varepsilon\right)$
$\leq\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{Q}^{\dagger}}(X)-1\right|\gtrsim\varepsilon\right)+\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\overline{\pi}_{\delta}}(X_{T-\delta})-1\right|\gtrsim\varepsilon\right).$
The inequality is due to the fact that $|ab-1|\geq\varepsilon$ implies $|a-1|\geq\frac{\varepsilon}{3}$ or $|b-1|\geq\frac{\varepsilon}{3}$ for $\varepsilon\in[0,1]$. It suffices to make both terms above $O(1)$. To bound the first term, we use the similar approach as in the proof of ˜18 in Thm.˜2:
@@
$\operatorname{Pr}_{X\sim\mathbb{Q}^{\dagger}}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\overline{\pi}_{\delta}}(X_{T-\delta})-1\right|\gtrsim\varepsilon\right)$ $\leq\operatorname{Pr}_{X\sim\mathbb{Q}}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\overline{\pi}_{\delta}}(X_{T-\delta})-1\right|\gtrsim\varepsilon\right)+\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\dagger})$
$\leq\overline{\pi}_{\delta}\left(\left|\frac{\mathrm{d}\pi}{\mathrm{d}\overline{\pi}_{\delta}}-1\right|\gtrsim\varepsilon\right)+\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\dagger})$
$\lesssim\frac{\operatorname{TV}(\overline{\pi}_{\delta},\pi)}{\varepsilon}+\varepsilon.$
Therefore, it suffices to make $\operatorname{TV}(\overline{\pi}_{\delta},\pi)\lesssim\varepsilon$. ∎